Use the following supply and demand equations.

Supply:p= 4 + 3q.

Demand:p= 2,132−q.

Use these equations to respond to the following questions.

(a) What is the market equilibrium?

(b) Under the market equilibrium, what is Total Surplus?

(c) Suppose the government enacts a price ceiling of ̄p= 2,000. What is ProducerSurplus, Consumer Surplus, Total Surplus, and Deadweight Loss?

(d) Instead, suppose that the government enacts a price ceiling of ̄p= 1,100. What is Producer Surplus, Consumer Surplus, Total Surplus, and Deadweight Loss?

(e) Under a price ceiling of ̄p= 2,000, what is the Shortage? What if ̄p= 1,100?
Answer #1

(a) The market equilibrium would be where the supply price would be equal to demand price, ie 4+39= 2132-g or 4q = 2128 or g 532 , and for p = 4 + 3q , we have p=4+3*532 or p= 1600 , which are the required equilibrium quantity and price.

(b) The graph is as below.

2 Supply 1500 1000 Demand 500 1000 1500 50 200

The total surplus is the area of the triangle ACE, which is TS-0.5 baseheight or TS = 0.5 * (2132-4) * 532 or TS= 5 60 IS dollars.

(c) In case of a price ceiling of p 2000 , the market equilibrium would not change. The reason being that the price ceiling of p 2000 would be non-binding or ineffective price ceiling, as the price ceiling of p 2000 means that more than $2000 price can not be charged in the market, and the market equilibrium is already lower than the price ceiling.

The producer surplus would be the area of triangle BCE, which is PS = 0.5*base*height or PS 0.5 * 532 * (1600-4 or PS 424536 dollars.

The consumer surplus would be the area of triangle ABE, which is CS = 0.5*base*height or CS = 0.5*532*(2132 - 1600) or CS = 141512 dollars.

The total surplus would be the area of PS and CS, ie TS = PS + CS or TS= 5 60 IS dollars, which is the same as before.

There is no dead-weight loss, since the price ceiling does not affect the market equilibrium.

(d) A price ceiling of \bar p = 1100 would be effective since it is below the market equilibrium. At this price, the quantity demanded is more than the quantity supplied. The quantity supplied would be as 1100 = 4 + 3q or q = 365.3333 , which would be the equilibrium quantity supplied, and traded. The graph is as below.

2 1500 Price Ceilin 1000 emand 500 1000 1500 200 500

The PS would be the area of triangle CBH, which is PS = 0.5*base*height or PS = 0.5*365.3333*(1100-4) or PS = 200202.6484 dollars.

The CS would be the area of region ABHG, which is the area of triangle AFG and area of rectangle FBHG, ie CS = 0.5*base*height + FB*BH or CS = 0.5*365.3333*(2132 - 1766.6667) + (1766.6667 - 1100)*365.3333 (since at point F, p=1766.6667) or CS = 310289.7556 dollars.

The total surplus would be the sum of PS and CS, ie TS = PS + CS or TS = 200202.6484 + 310289.7556 or TS = 510492.404 dollars.

The dead-weight loss would be the difference between the old TS and the new TS, which is DWL = 566048 - 510492.404 = 55555.596 dollars, which is also the area of triangle GHE.

(e) Under the price ceiling of p 2000 , there is no shortage since the price ceiling is above the market equilibrium.

If price ceiling is \bar p = 1100 , the quantity supplied is q = 365.3333 , while the quantity demanded would be as 1100 = 2132 - q or q = 1032 . The shortage in this case would hence be Shortage = 1032 - 365.3333 = 666.6667 units, which is the demand not met by the supply due to the price ceiling.